11. Let the correlation coefficient of X and Y be ρ(X,Y)-N C XY VVar(X)VVar(Y) -p(X, Y). (Y) Show that ρ(-3X,-2Y-a(X, Y) and ρ(X-2Y)
Answer #1

Ans. Since variances as well as coavariances are independent of change of origin but depend on change of scale.

The correlation coefficient is ratio of covariance and variance, being independent of change of origin and sclae but depend on sign.

Conn (x,y) xyCov(x,y) = Van (X) Von (Y) Now Van (-3x) (-3)2 van (x) -g vorn (X) = -2)-Van ( Cov(-3x,-2y) Cov(xN) Cos (X1)Now Van (-2y) 4 van (y) = Cov (x,-2Y) Val -J(x,y)